i need help in solving with a coin and intrest problem> it goes like this an invester has 500$ more invested at 7% than he does at 5%. if his annual interest is $515, how much does he have invested at each rate?
Found 2 solutions by wuwei96815, stanbon:
Answer by wuwei96815(245)

If we let x = amount invested at 5% and x+500 = the amount invested at 7%, we can write and solve the following equation:
5%x + 7%(x+500) = $515.

5%x + 7%(x+500) = $515
0.05x + 0.07x + 35 = 515
0.12x = 515 - 35 = 480
x = 480/0.12 = 4000


x = $4000
x+500 = $4500

Answer by stanbon(57967)

Let amount invested at 5% be "x".
The amount invested at 7% is "x+500"
EQUATION:
Interest + Interest= $515
0.05x + 0.07(x+500)=515
Multiply through be 100 to get:
5x+7x+3500=51500
12x=48000
x=4000 (amount invested at 5%
x+500=4500 (amount invested at 7%)
Cheers,
Stan H.